solve by the elimination method
4a+3b=-9
-4a+b=13

Answer by jim_thompson5910(35256)

Hint:

Add the equations:

4a+3b=-9
-4a+b=13
---------
0a+4b=4

4b = 4

b = 4/4

b = 1

Now use this value of 'b' to find the value of 'a'